OpenStudy (anonymous):

What is the equation in point−slope form of the line passing through (−3, −4) and (0, 2)? (y − 8) = 2(x − 3) (y + 3) = −2(x + 4) (y − 2) = −2(x + 4) (y + 4) = 2(x + 3)

OpenStudy (anonymous):

@AlexandervonHumboldt2 @Compassionate @Callisto @phi @mathstudent55 @iGreen

OpenStudy (igreen):

First, plug the points into the slope formula to find the slope. \(\sf m = \dfrac{y_2-y_1}{x_2-x_1}\) \(\sf m = \dfrac{2+4}{0+3}\)

OpenStudy (anonymous):

(y + 4) = 2(x + 3)?